# UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction

### v2

@@ A.1 Link Prediction Methods for Triple-based KGs
Most existing techniques in KG representation learning are proposed for triple-based KGs. Among these techniques, knowledge graph embedding (KGE) models Bordes et al. (2013); Sun et al. (2018) have received extensive attention due to their effectiveness and simplicity. The idea is to project entities and relations in the KG to low-dimensional vector spaces, utilizing KGE scoring functions to measure the plausibility of triples in the embedding space. Typical methods include TransE Bordes et al. (2013), RotatE Sun et al. (2018), and ConvE Dettmers et al. (2018).
@@
Depending on the KGE model alone has limitation of capturing complex graph structures, whereas augmenting global structural information with a graph neural network (GNN) Vashishth et al. (2019); Nathani et al. (2019); Xu et al. (2023b) proves to be an effective approach for enhancement. The paradigm of combining GNN as encoder with KGE scoring function as decoder helps to enhance the performance of KGE scoring function. These GNN methods design elaborate message passing mechanisms to capture the global structural features. Typically, CompGCN Vashishth et al. (2019) aggregates the joint embedding of entities and relations in the neighborhood via a parameter-efficient way and MA-GNN Xu et al. (2023b) learns global-local structural information based on multi-attention. These methods achieve impressive results on triple-based KGs but are hard to generalize to beyond-triple KGs.
Appendix B Results on JF17K
Table 6 shows the experimental results on JF17K. Due to the absence of a validation set in the JF17K dataset and the different ways of dividing the dataset across various baselines, we adopt the results reported in the original paper. Consistent with previous experiments on hyper-relational knowledge graphs, we also achieve state-of-the-art performance on JF17K among all baselines. In particular, we achieved 1.7 (2.9%) points in MRR and 1.5 (2.9%) points in Hits@1 compared to the method StarE which also utilises a graph neural network encoding and a simple transformer decoding, indicating that our hierarchical GNN HiSL could better capture the structure of hyper-relational facts.
Table 6: Link prediction on JF17K. All results of baselines are taken from the original paper. Best results are in bold.
Model JF17K
subject/object all entities
MRR Hits@1 Hits@10 MRR Hits@1 Hits@10
m-TransH 0.206 0.206 0.462 - - -
NaLP 0.221 0.165 0.331 0.366 0.290 0.516
HINGE 0.431 0.342 0.611 0.517 0.436 0.675
NeuInfer 0.449 0.361 0.624 0.473 0.397 0.618
StarE 0.574 0.496 0.725 - - -
HyTransformer 0.582 0.501 0.742 - - -
UniHR 0.591 0.511 0.745 0.621 0.545 0.768
Appendix C Decoder Analysis
To explore the effectiveness of our UniHR encoding further, we pair UniHR with different decoders and evaluated them on triple prediction task. In addition to the previously mentioned unified framework UniHR + Transformer, we also experiment on UniHR + ConvE with two scoring strategies. The ConvE Dettmers et al. (2018) is the decoder customized for triples and its scoring function is $vec\left(\sigma\left(\left[\mathbf{\tilde{h}}_{h};\mathbf{\tilde{e}}_{r}\right]*\psi\right)\right)$, where $\mathbf{\tilde{h}}_{h}$ and $\mathbf{\tilde{e}}_{r}$ represent reshaped 2D embeddings of head entity $h$ and relation $r$, and * is a convolution operator. The $vec\left(\cdot\right)$ and $\psi$ are denoted as the vectorization function and a set of convolution kernels.
Table 8 shows the results of triple prediction on three benchmark datasets. Among all baselines, Quate, Bique, Neural-LP, Drum, and AnyBURL struggle to model the mapping relationship between atomic facts and nested facts. Furthermore, prior works Chung and Whang (2023) do not guarantee that all atomic facts in the nested fact test set are present in the training set as entities, which shifts the problem from a transductive setting to an inductive setting, leading to significant performance gaps between these baselines. On most metrics, our method outperforms BiVE and NestE which are specifically modeled for nested facts. Notably, NestE fully preserves the semantics of atomic facts. However, on the FBHE dataset, UniHR + ConvE achieves an improvement of 0.58 (6.4%) points in MRR and 0.24 (2.4%) points in Hits@10 compared to the state-of-the-art model NestE and the second-best performance after UniHR + Transformer on the FBH and DBHE datasets, demonstrating UniHR’s powerful graph structure encoding capabilities. We also carry out ablation experiments on UniHR + ConvE as shown in Table 8. Performance declines are observed after removing any part of the HiSL module, showing the significance of HiSL for hierarchical encoding.
@@ Appendix C Decoder Analysis
Table 7: Base link prediction on FBHE, FBH and DBHE. All baselines’ results are taken from Xiong et al. (2024). The best results among all models are written bold, while the second are underlined. The $s_{f}$ and $s_{h}$ denote $\left(f,has\;head\;entity,h\right)\ \left(f,has\;tail\;entity,t\right)$ and $\left(h,r,t\right)$ two types of scoring method respectively. For BiVE Chung and Whang (2023) and NestE Xiong et al. (2024), we pick their variants with best performance.
@@
UniHR + Transformer 0.401 0.619 0.300 0.455
Table 8: Triple prediction on FBHE, FBH and DBHE. All baselines’ results are taken from Xiong et al. (2024). The best results among all models are written bold. For BiVE Chung and Whang (2023) and NestE Xiong et al. (2024), we pick their variants with best performance.
@@
QuatE Zhang et al. (2019) 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
BiQUE Guo and Kok (2021) 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
Neural-LP Yang et al. (2017) 115016.6 0.070 0.073 90000.4 0.238 0.274 21130.5 0.170 0.209
DRUM Sadeghian et al. (2019) 115016.6 0.069 0.073 90000.3 0.261 0.274 21130.5 0.166 0.209
AnyBURL Meilicke et al. (2019) 108079.6 0.096 0.108 83136.8 0.191 0.252 20530.8 0.177 0.214
BiVE Chung and Whang (2023) 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
NestE Xiong et al. (2024) 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
@@
Table 9: The statistics of diverse knowledge graphs dataset, where “with Q(%)" and “Arity" column respectively denote the number of facts with auxiliary key-value pairs and the range of arity of hyper-relational facts, the “with AF(%)" column denotes the number of atomic facts in nested facts.
Dataset Atomic Fact Entities Relations Train Valid Test with Q(%) Arity Nested Fact Nested Relation with AF(%) Period
Hyper-relational Knowledge Graph
WikiPeople 369866 34825 178 294439 37715 37712 9482(2.6%) 2-7 - - - -
WD50K 236507 47155 531 166435 23913 46159 32167(13.6%) 2-67 - - - -
JF17K 100947 28645 501 76379 - 24568 46320(45.9%) 2-6 - - - -
Nested Factual Knowledge Graph
FBH 310116 14541 237 248094 31011 31011 - - 27062 6 33157 -
FBHE 310116 14541 237 248094 31011 31011 - - 34941 10 33719 -
DBHE 68296 12440 87 54636 6830 6830 - - 6717 8 8206 -
Temporal Knowledge Graph
wikidata12k 40621 12554 24 32497 4062 4062 - - - - - 19-2020
Multiple types of Knowledge Graph
wikimix 409566 43832 184 326936 41777 3525(TKG)/37328(HKG) 9098(2.2%) 2-7 - - - 19-2020
@@
Algorithm 1 Message passing process of HiSL
Input: $\mathcal{G}^{HiDR}=\{\mathcal{V}_{a}\cup\mathcal{V}_{r}\cup\mathcal{V}_{f},\mathcal{R}_{a}\cup\mathcal{R}_{n}\cup\mathcal{R}_{c},\mathcal{F}_{a}\cup\mathcal{F}_{c}\cup\mathcal{F}_{n}\}$; the number of encoder layers $L$.
Output: The node embedding matrix $\mathbf{H}^{L}$; the edge embedding matrix $\mathbf{E}^{L}$.
1:  Initialize the embedding matrix of atomic nodes $\mathbf{H}_{a}^{0}\in\mathbb{R}^{\left|\mathcal{V}_{a}\right|\times d}$.
2:  Initialize the embedding matrix of three types of relations $\mathbf{E}^{0}=\{\mathbf{E}_{a}^{0},\mathbf{E}_{c}^{0},\mathbf{E}_{n}^{0}\}\in\mathbb{R}^{\left|\mathcal{R}_{a}\cup\mathcal{R}_{c}\cup\mathcal{R}_{n}\right|\times d}$.
3:  Initialize the embedding matrix of relation nodes $\mathbf{H}_{r}^{0}\leftarrow\mathbf{E}_{a}^{0}\cdot\mathbf{W}_{r}\in\mathbb{R}^{\left|\mathcal{V}_{r}\right|\times d}$.
4:  for  $h,r,t\in\mathcal{F}_{a}$ do
5:     Initialize the embedding of fact nodes $\mathbf{h}_{f}^{0}\leftarrow f_{m}([\mathbf{h}_{h};\mathbf{h}_{r};\mathbf{h}_{t}])$. $\triangleright$ Representation Initialization
6:  end for
7:  for  $l\leftarrow 1$ to $L$ do
8:     for $i\in\mathcal{V}$ do
9:        $\mathcal{M}^{l}_{i}\leftarrow\mathbf{h}_{i}^{l-1}$. $\triangleright$ Intra-fact Message Passing
10:        for  $i,r,j\in\mathcal{F}_{c}$ do
11:           $\alpha_{i,j}^{l-1}\leftarrow\frac{\exp\left(\mathbf{W}^{l-1}\left(\sigma\left(\mathbf{W}_{in}^{l-1}\mathbf{h}_{i}^{l-1}+\mathbf{W}_{out}^{l-1}\mathbf{h}_{j}^{l-1}\right)\right)\right)}{\sum_{j^{\prime}\in\mathcal{N}_{i}}{\exp\left(\mathbf{W}^{l-1}\left(\sigma\left(\mathbf{W}_{in}^{l-1}\mathbf{h}_{i}^{l-1}+\mathbf{W}_{out}^{l-1}\mathbf{h}_{j^{\prime}}^{l-1}\right)\right)\right)}}$.
12:           $\mathcal{M}^{l}_{i}\leftarrow\{\alpha_{i,j}^{l-1}\cdot\mathbf{W}_{out}^{l-1}\mathbf{h}_{j}^{l-1}\}\cup\mathcal{M}^{l}_{i}$.
13:        end for
14:     end for
15:     for $i\in\mathcal{V}$ do
[…]
17:     end $\triangleright$ Inter-fact Message Passing
18:     for $i\in\mathcal{V}$ do
19:        $\mathcal{M}^{l}_{i}\leftarrow\{\mathbf{W}_{self}^{{l-1}}\mathbf{h}_{i}^{{l-1}}\}$.
20:        for  $r,j\in\mathcal{N}_{i}$ do
21:           $\mathbf{m}^{l-1}_{(i,r,j)}\leftarrow\sigma\left(\omega_{\tau\left(r\right)}^{l-1}\right)\mathbf{W}_{{\lambda}\left(r\right)}^{{l-1}}\phi\left(\mathbf{h}_{{j}}^{{l-1}},\mathbf{e}_{{r}}^{{l-1}}\right)$.
22:           $\mathcal{M}^{l}_{i}\leftarrow\{\mathbf{m}^{l-1}_{(i,r,j)}\}\cup\mathcal{M}^{l}_{i}$.
23:        end for
24:     end for
25:     for $i\in\mathcal{V}$ do
[…]
27:     end for
28:     $\mathbf{E}^{l}\leftarrow\mathbf{W}_{rel}^{l-1}\mathbf{E}^{l-1}$.
29:  end for
30:  return  $\mathbf{H}^{L},\mathbf{E}^{L}$
Appendix D Datasets Statistics
Table 9 shows the details of the three hyper-relational knowledge graph benchmark datasets: WikiPeople, WD50K, JF17K, three nested factual knowledge graph benchmark datasets: FBH, FBHE, DBHE,and the temporal knowledge graph benchmark dataset wikidata12k. Among them, WikiPeople is a dataset derived from Wikidata Vrandečić and Krötzsch (2014) concerning entities type “human”. WikiPeople filter out the elements which have at least 30 mentions as key-value pairs. WD50K is a high-quality dataset extracting from Wikidata statements and avoiding the potential data leakage which allows triple-based models to memorize main fact in the H-Facts of test set. The “with Q(%)" column in Table 9 denote the number of facts with auxiliary key-value pairs and the “Arity" column denote range of the number of entities in hyper-relational facts. The nested factual knowledge graph datasets FBH and FBHE Chung and Whang (2023) are constructed based on FB15k237 Toutanova and Chen (2015) from Freebase Bollacker et al. (2008) while DBHE is based on DB15K Liu et al. (2019) from DBpedia Lehmann et al. (2015). FBH contains nested facts that can be only inferred inside the atomic facts, while FBHE and DBHE contain externally-sourced nested relation crawling from Wikipedia articles, e.g., NextAlmaMater and SucceededBy. Temporal knowledge graph dataset wikidata12K is also a subset of Wikidata Vrandečić and Krötzsch (2014), which represents the time information $\tau\in\mathcal{T}$ as time intervals.
Appendix E Pseudo-code of HiSL
The pseudo code of HiSL is outlined in Algorithm 1.
Appendix F Hyperparameter Settings

### v3

@@ 0.A.1 Link Prediction on Triple-based KGs
Most existing techniques in KG representation learning are proposed for triple-based KGs. Among these techniques, knowledge graph embedding (KGE) models [4, 29] have received extensive attention due to their effectiveness and simplicity. The idea is to project entities and relations in the KG to low-dimensional vector spaces, utilizing KGE scoring functions to measure the plausibility of triples in the embedding space. Typical methods include TransE [4], RotatE [29], and ConvE [9].
@@
Depending on the KGE model alone has limitation of capturing complex graph structures, whereas augmenting global structural information with a graph neural network (GNN) [31, 25, 41] proves to be an effective approach for enhancement. The paradigm of combining GNN as encoder with KGE scoring function as decoder helps to enhance the performance of KGE scoring function. These GNN methods design elaborate message passing mechanisms to capture the global structural features. Typically, CompGCN [31] aggregates the joint embedding of entities and relations in the neighborhood via a parameter-efficient way and MA-GNN [41] learns global-local structural information based on multi-attention. These methods achieve impressive results on triple-based KGs but are hard to generalize to beyond-triple KGs.
@@
Table 10: Base link prediction on FBHE, FBH and DBHE. All baselines’ results are taken from [37]. The best results among all models are written bold, while the second are underlined. The $s_{f}$ and $s_{h}$ denote $\left(f,has\;head\;entity,h\right)\ \left(f,has\;tail\;entity,t\right)$ and $\left(h,r,t\right)$ two types of scoring method respectively. For BiVE [7] and NestE [37], we pick their variants with best performance.
@@
UniHR + Transformer 0.401 0.619 0.296 0.448
Appendix 0.B Decoder Analysis
To explore the effectiveness of our UniHR encoding further, we pair UniHR with different decoders and evaluated them on triple prediction task. In addition to the previously mentioned unified framework UniHR + Transformer, we also experiment on UniHR + ConvE with two scoring strategies. The ConvE [9] is the decoder customized for triples and its scoring function is $vec\left(\sigma\left(\left[\mathbf{\tilde{h}}_{h};\mathbf{\tilde{e}}_{r}\right]*\psi\right)\right)$, where $\mathbf{\tilde{h}}_{h}$ and $\mathbf{\tilde{e}}_{r}$ represent reshaped 2D embeddings of head entity $h$ and relation $r$, and * is a convolution operator. The $vec\left(\cdot\right)$ and $\psi$ are denoted as the vectorization function and a set of convolution kernels.
Table 11: Triple prediction on FBHE, FBH and DBHE. All baselines’ results are taken from [37]. The best results among all models are written bold. For BiVE [7] and NestE [37], we pick their variants with best performance.
@@ Appendix 0.B Decoder Analysis
QuatE [45] 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
BiQUE [16] 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
Neural-LP [42] 115016.6 0.070 0.073 90000.4 0.238 0.274 21130.5 0.170 0.209
DRUM [28] 115016.6 0.069 0.073 90000.3 0.261 0.274 21130.5 0.166 0.209
AnyBURL [24] 108079.6 0.096 0.108 83136.8 0.191 0.252 20530.8 0.177 0.214
BiVE [7] 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
NestE [37] 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
@@
Table 11 shows the results of triple prediction on three benchmark datasets. Among all baselines, Quate, Bique, Neural-LP, Drum, and AnyBURL struggle to model the mapping relationship between atomic facts and nested facts. Furthermore, prior works [7] do not guarantee that all atomic facts in the nested fact test set are present in the training set as entities, which shifts the problem from a transductive setting to an inductive setting, leading to significant performance gaps between these baselines. On most metrics, our method outperforms BiVE and NestE which are specifically modeled for nested facts. Notably, NestE fully preserves the semantics of atomic facts. However, on the FBHE dataset, UniHR + ConvE achieves an improvement of 0.58 (6.4%) points in MRR and 0.24 (2.4%) points in Hits@10 compared to the state-of-the-art model NestE and the second-best performance after UniHR + Transformer on the FBH and DBHE datasets, demonstrating UniHR’s powerful graph structure encoding capabilities. We also carry out ablation experiments on UniHR + ConvE as shown in Table 11. Performance declines are observed after removing any part of the HiSL module, showing the significance of HiSL for hierarchical encoding.
@@
Appendix 0.C Hyperparameter Settings